OpenStudy (anonymous):

Lane has a gift card worth $300 for a local entertainment store. Movies cost $20 each and newly released video games cost $50 each. In order to use the gift card, Lane must purchase at least seven items. Which is a combination of movies and newly released video games that Lane could purchase using his gift card? help me please

OpenStudy (igreen):

We can use a system of equations. First equation: \(20x + 50y = 300\) Where x = movies, and y = video games. Second equation: \(x + y = 7\). This is for the combination.

OpenStudy (igreen):

Now we just need to solve them.

OpenStudy (anonymous):

so its After 7 weeks, Randall is $225 in debt.

OpenStudy (igreen):

What are you talking about..?

OpenStudy (anonymous):

sorry I'm out of subject so is it going to be 6 movies and 5 video games

OpenStudy (igreen):

No the total has to be 7. 6 + 5 = 11.

OpenStudy (anonymous):

2 movies and 4 video games how 11

OpenStudy (igreen):

You said 6 movies and 5 video games.

OpenStudy (igreen):

2 movies and 4 video games is 6.

OpenStudy (anonymous):

6 movies and 5 video games is the answer :)

OpenStudy (igreen):

If you graph it you get \(x = \dfrac{1}{6}\), \(y = 5 \dfrac{1}{3}\).

OpenStudy (anonymous):

am i wrong i'm not good at graph

OpenStudy (igreen):

Your answer is 1/6 movie and 5 1/3 video game.

OpenStudy (anonymous):

thanks

OpenStudy (igreen):

@barbie15